Given an underground total mining cost of $8.50 per tonne, surface mining costs of $2.25 per...

  1. Given an underground total mining cost of $8.50 per tonne, surface mining costs of $2.25 per tonne of ore and surface mining stripping costs of $1.95 per tonne of waste. Calculate the break even stripping ratio, (i.e with zero profit) between surface and underground mining. That is, the point at which underground mining should be contemplated.   (5 points)

Please type your answer, Please do not write on the paper and post it
Answer #1

Break even stripping ratio = (UG - OP) /SG

UG - underground mining cost per tonn = $8.5

OP- Surface mining cost per tonn of ore = $2.25

SG - surface mining stripping cost per tonn of waste = $1.95

BREAK EVEN STRIPPING RATIO at zero profit,

= (8.5 - 2.25)/1.95 = 3.205

Ans: 3.205